Find the foci of this hyperbola by using the equation (x+9)^2/1 - (x+5)^2/49 = 1.Can someone please help me.I can seem to figure it out.
Answer by lwsshak3(11628)

Find the foci of this hyperbola by using the equation (x+9)^2/1 - (x+5)^2/49 = 1
..
standard form of hyperbola:(x-h)^2/a^2-(y-k)^2/b^2=1
given:(x+9)^2/1-(x+5)^2/49 = 1
I will assume this was submitted in error, that is,(x+5) of the second term should instead be (y+5)as:(x+9)^2/1-(y+5)^2/49 = 1
If this is the case, this is hyperbola centered at (-9,-5) with horizontal transverse axis (opens sideways)
a^2=1
a=1
b^2=49
b=7
length of foci,c, =sqrt(a^2+b^2)=sqrt(50)=7.07
To find the coordinates of the foci, you know they lie somewhere on the transverse axis,y=-5, so the y-coordinate=-5
The x-coordinate=+-from the center x-coordinate, that is, -9=-7.07=-16.07 and -2.07
ans:
The foci is located at (-16.07,-5) and (-2.07,-5)